Please help me solve this equation: nCr-3+3(nCr-2)+3(nCr-1)+nCr
Found 3 solutions by ikleyn, greenestamps, Alan3354:
Answer by ikleyn(52792)

.

There is no ANY equation in your post.



Answer by greenestamps(13200)



Empirically (by looking at examples) the given expression is equivalent to (n+3)Cr. e.g....

6C3+3*6C4+3*6C5+6C6 = 20+3(15)+3(6)+1 = 84 = 9C6
5C1+3*5C2+3*5C3+5C4 = 5+3(10)+3(10)+5 = 70 = 8C4
7C2+3*7C3+3*7C4+7C5 = 21+3(35)+3(35)+21 = 252 = 10C5

Proving it algebraically looks like a nightmare, unless there is some trick I don't see.

Answer by Alan3354(69443)

That's not an equation.
Equations have equal signs.